# Danish Fungi 2024 DF20 Danish Fungi 2020

### v2

@@ 3 FungiTastic Benchmarks
The diversity and unique features of the FungiTastic dataset allow for the evaluation of various fundamental computer vision and machine learning problems. We present several distinct benchmarks, each with its own evaluation protocol. This section provides a detailed description of each challenge and the corresponding evaluation metrics. Metrics are further defined in Appendix A.
@@
Closed-set classification: The FungiTastic dataset is a challenging dataset with many visually similar species, heavy long-tailed distribution, and considerable distribution shifts over time. Since the fine-grained close-set classification methodology is well-defined, we follow the widely accepted standard, and we, apart from accuracy, use the macro-averaged F1-score ($\text{F}_{1}^{m}$).
@@
Open-set classification: In the Atlas of Danish Fungi (our data source), new species are continuously added to the database, including previously unreported species. This ongoing data acquisition enables a yearly data split with a natural class distribution shift (see Figure 5), and many species in the test data are absent in the training set. We follow a widely accepted methodology, and we propose an AUC as the main metric. Besides, we calculate $\text{F}_{1}^{m}$ for two classes: "unknown" and "known."
@@
Few-shot classification: All the categories with less than five samples in the FungiTastic, usually uncommon and rare species, form the few-shot subset. Being capable of recognizing those is of high interest to the expert. Since the few-shot dataset has no severe class imbalance like the other FungiTastic subsets, this benchmark’s main metric is top-1 accuracy. The F-1 score and top-k total accuracy are also reported. This challenge does not have any “unknown” category.
@@
Chronological classification: Each observation in the FungiTastic (FungiTastic) dataset has a timestamp, allowing the study of species distribution changes over time. Fungi distribution is seasonal and influenced by weather, such as recent precipitation. New locations may be added over time, providing a real-world benchmark for domain adaptation methods, including online, continual, and test-time adaptation. The test dataset consists of fungi images ordered chronologically, meaning a model processing an observation at time $t$ can access all observations with timestamps $t^{\prime}<t$.
@@
Classification beyond 0–1 loss function: Evaluation of classification networks is typically based on the 0-1 loss function, such as the mean accuracy, which applies to the metrics defined for the previous challenges as well. In practice, this often falls short of the desired metric since not all errors are equal. In this challenge, we define two practical scenarios: In the first scenario, confusing a poisonous species for an edible one (false positive edible mushroom) incurs a much higher cost than that of a false positive poisonous mushroom prediction. In the second scenario, the cost of not recognizing that an image belongs to a new species should be higher.
@@
Segmentation: While acquiring human-annotated ground-truth segmentation masks can be resource-intensive, segmentation plays a critical role in advanced recognition methods and evaluation protocols for fine-grained image classification [5, 43]. High-accuracy segmentation of fungal images not only supports such methods but also enables automated analyses, such as assessing relationships between species-specific morphological features and environmental factors. This facilitates deeper insights into ecological and morphological patterns across locations, enhancing both model robustness and biological relevance. FungiTastic is suitable for evaluating semantic segmentation with the standard mean intersection over union (mIoU) metric and instance segmentation with the average precision (AP) metric.
Figure 5: Class distribution shift on the FungiTastic–M dataset.The long-term data acquisition captures a phenomenon related to natural changes in species presence, i.e., class prior shift. Sorted in descending order based on their occurrence in the training set (color-coded to indicate relative frequencies in the training, validation, and test sets). The training set includes data from before 2022 (215 species), the validation set from 2022 (196 species), and the test set from 2023 (193 species).
@@ 4 Baseline Experiments
In this section, we provide a variety of weak and strong baselines based on state-of-the-art architectures and methods for three FungiTastic benchmarks. Bellow, we report results only for the closed-set, few-shot learning, and zero-shot segmentation, but other baselines will be provided later in the supplementary materials, in the documentation, or on the dataset website.
@@ 4.1 Closed-set image classification
We train a variety of state-of-the-art CNN architectures to establish some baselines for closed-set classification on the FungiTastic and FungiTastic–M. All selected architectures were optimized with Stochastic Gradient Descent with momentum set to 0.9, SeeSaw loss [54], a mini-batch size of 64, and Random Augment [8] with a magnitude of 0.2. The initial learning rate was set to 0.01 (except for ResNet and ResNeXt, where we used LR=0.1), and it was scheduled based on validation loss.
Results: Similarly to other fine-grained benchmarks, while the number of params, complexity of the model, and training time are more or less the same, the transformer-based architectures achieved considerably better performance on both FungiTastic and FungiTastic–M and two different input sizes (see Table 3). The best-performing model, BEiT-Base/p16, achieved $\text{F}_{1}^{m}$ just around 40%, which shows the severe difficulty of the proposed benchmark.
Table 3: Closed-set fine-grained classification FungiTastic and FungiTastic–M. A set of selected state-of-the-art CNN- (top section) and Transformer-based (bottom section) architectures evaluated on the FungiTastic and FungiTastic–M test sets. All reported metrics show the challenging nature of the dataset. The best result for each metric is highlighted.
FungiTastic–M – 2242 FungiTastic – 2242 FungiTastic–M – 3842 FungiTastic – 3842
Architectures Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$
ResNet-50 61.7 79.3 35.2 62.4 77.3 32.8 66.3 82.9 39.8 66.9 80.9 36.3
ResNeXt-50 62.3 79.6 36.0 63.6 78.3 33.8 67.0 84.0 39.9 68.1 81.9 37.5
EfficientNet-B3 61.9 79.2 36.0 64.8 79.4 34.7 67.4 82.8 40.5 68.2 81.9 37.2
EfficientNet-v2-B3 65.5 82.1 38.1 66.0 80.0 36.0 70.3 85.8 43.9 72.0 84.7 41.0
ConvNeXt-Base 66.9 84.0 41.0 67.1 81.3 36.4 70.2 85.7 43.9 70.7 83.8 39.6
ViT-Base/p16 68.0 84.9 39.9 69.7 82.8 38.6 73.9 87.8 46.3 74.9 86.3 43.9
Swin-Base/p4w12 69.2 85.0 42.2 69.3 82.5 38.2 72.9 87.0 47.1 74.3 86.3 43.3
BEiT-Base/p16 69.1 84.6 42.3 70.2 83.2 39.8 74.8 88.3 48.5 75.3 86.7 44.5
@@ 4.2 Few-shot image classification
Three baseline methods are implemented. The first baseline is standard classifier training with the Cross-Entropy (CE) loss. The other two baselines are nearest-neighbor classification and centroid prototype classification based on deep image embeddings extracted from large-scale pre trained vision models, namely CLIP [41], BioCLIP [47] and DINOv2 [33].
@@
Standard deep classifiers are trained with the CE loss to output the class probabilities for each input sample. Nearest neighbors classification ($\mathbf{k}$-NN) constructs a database of training image embeddings. At test time, $k$ nearest neighbors are retrieved, and the classification decision is made based on the majority class of the nearest neighbors. Nearest-centroid-prototype classification constructs a prototype embedding for each class by aggregating the training data embeddings of the given class. The classification is performed based on the image embedding similarity to the class prototypes. These methods are inspired by prototype networks proposed in [45].
@@
Results: While DINOv2 [33] embeddings greatly outperform CLIP [41] embeddings, BioCLIP [47] (CLIP finetuned on biological data) outperforms them both, highlighting the dominance of domain-specific models. Further, the centroid-prototype classification always outperforms the nearest-neighbor methods in terms of accuracy, while the nearest-neighbor wins over the centroid-prototype in F-score. Finally, the best standard classification models trained on the in-domain few-shot dataset underperform both DINOv2 and BioCLIP embeddings in the F-score, which shows the power of methods tailored to the few-shot setup. For results summary, refer to Table 4.
@@
Table 4: Few shot classification on FungiTastic–Few-shot. (Left) – Pretrained deep descriptors with the nearest centroid and 1-NN classification. All pre-trained models are based on the ViT-B architecture, CLIP, and BioCLIP with patch size 32 and DINOv2 with patch size 16. (Right) – Standard classification with cross-entropy-loss. Evaluated on the FungiTastic–Few-shot test set.
[…]
CLIP 1-NN 6.1 2.8 4.2 5.3 9.9 11.5
centroid 7.2 2.2 13.0 3.6 7.4 12.5 12.5
DINOv2 1-NN 17.4 8.4 11.9 18.0 28.6 21.4
centroid 17.9 5.9 27.8 12.2 17.8 23.4 31.6
BioCLIP 1-NN 18.8 9.1 12.8 21.3 30.0 29.4
centroid 21.8 6.8 32.6 12.7 23.5 32.9 40.3
Architecture Input Top1   $\text{F}_{1}^{m}$ Top3
BEiT-B/p16 224$\times$224 11.0 2.1 17.4
384$\times$384 11.4 2.1 18.4
ConvNeXt-B 224$\times$224 14.0 2.7 23.1
384$\times$384 15.4 2.9 23.6
ViT-B/p16 224$\times$224 13.9 2.7 21.5
384$\times$384 19.5 3.7 29.0
4.3 Experiments with additional data
@@
We provide baseline experiments using tabular metadata (habitat, month, substrate) based on our previous work [40]. The results in Table LABEL:tab:exps:tab_data show that all the attributes improve all the metrics. Individually, the addition of the habitat attribute results in the biggest gains (+2.27% of accuracy), followed by substrate (+1.15%) and month (+0.87%). Overall, habitat was the most efficient way to improve performance. With the combination of Habitat, Substrate, and Month, we improved the EfficientNet-B3 model’s performance on FungiTastic-Mini by 3.62%, 3.42% and 7.46% in Top1, Top3, and F1, respectively, indicating the gains are mostly orthogonal. Using the MetaSubstrate instead of Substrate resulted in performance lower by 0.2%, 0.5%, and 0.3% in Top1, Top3, and F1, respectively.
@@
Table 5: Ablation on a combination of selected observation-related metadata. Utilizing a simple yet effective approach based on our previous work [40], we measure performance improvement using Habitat, Substrate, and Month and their combination. We also test how replacing Substrate with 31 variables for MetaSubstrate with nine variables with better interoperability to non-experts influences the performance. Evaluated with EfficientNet-B3 on FungiTastic-Mini test set.
@@ 4.4 Segmentation
Results: The baseline method achieved an average per-image IoU of 89.36%. While the model exhibits strong zero-shot performance, it sometimes fails to detect mushrooms. These instances often involve small mushrooms, where a higher input resolution could enhance detection, and atypical mushrooms, such as very thin ones. Another common issue is SAM’s tendency to miss mushroom stems. The results for the simplified foreground-background segmentation task (rather than instance-level semantic segmentation) underline the need for further development of domain-specific models. Qualitative results, including random images and examples where the segmentation performs best and worst, are reported in Figures 8, 8, and 8.
@@
Figure 6: Foreground-background segmentations on randomly selected FungiTastic–M photos. Highlighted pixels correspond to: true positives, false positives, and false negatives.
Figure 7: Foreground-background segmentations on FungiTastic–M with the highest IoU. Highlighted pixels correspond to: true positives.
Figure 8: Foreground-background segmentations on FungiTastic–M with the lowest IoU value. Highlighted pixels correspond to: false positives and false negatives.

### v3

@@ 4 FungiTastic Benchmarks
The diversity and unique features of the FungiTastic dataset allow for the evaluation of various fundamental computer vision and machine learning problems. We present several benchmarks, each with its own evaluation protocol. This section provides a detailed description of each challenge and the corresponding evaluation metrics. Metrics are further defined in Appendix A.
@@
Closed-set classification: The FungiTastic dataset is a challenging dataset with many visually similar species, a heavy long-tailed distribution, and considerable distribution shifts over time. Since the fine-grained closed-set classification methodology is well-defined, we follow the widely accepted standard, and we, apart from accuracy, use the macro-averaged F1-score ($\text{F}_{1}^{m}$).
@@
Open-set classification: In the Atlas of Danish Fungi (our data source), new species are continuously added to the database, including previously unreported species. This long-term ongoing data acquisition enables a yearly data split with a natural class distribution shift (see Figure 5), and many species in the test data are absent in the training set. We follow a widely accepted methodology, and we propose to use an AUC as the main metric. Besides, we calculate True Negative Rate at 95% True Positive Rate (TNR95) metric.
@@
Few-shot classification: All the categories with less than five samples, usually uncommon and rare species, form the few-shot subset. Being capable of recognizing those is of high interest to the experts. Since the few-shot dataset has no severe class imbalance like the other FungiTastic subsets, the main metric is Top1 accuracy. The macro-averaged F1-score ($\text{F}_{1}^{m}$) and Top3 total accuracy are also reported. This challenge does not have any “unknown” category.
@@
Chronological classification: Each observation in the FungiTastic dataset has a timestamp, allowing the study of species distribution changes over time. Fungi distribution is seasonal and influenced by weather, such as recent precipitation. New locations may be added over time, providing a real-world benchmark for domain adaptation methods, including online, continual, and test-time adaptation. The test dataset consists of fungi images ordered chronologically, meaning a model processing an observation at time $t$ can access all observations with timestamps $t^{\prime}<t$.
@@
Classification beyond 0–1 loss function: Evaluation of classification networks is typically based on the 0–1 loss function, such as the mean accuracy, which also applies to the metrics defined for the previous challenges. This often falls short of the desired metric in practice since not all errors are equal. In this challenge, we define two practical scenarios: In the first scenario, confusing a poisonous species for an edible one (false positive edible mushroom) incurs a much higher cost than that of a false positive poisonous mushroom prediction. In the second scenario, the cost of not recognizing that an image belongs to a new species should be higher.
@@
Segmentation: Acquiring human-annotated segmentation masks can be resource-intensive, yet segmentation is vital for advanced recognition and fine-grained classification methods [8, 53]. Accurate segmentation of fungal images supports these methods and enables automated analysis of species-specific morphological and environmental relationships and revelation of ecological and morphological patterns across locations. With its annotations, FungiTastic-M is built to accommodate semantic segmentation using the standard mean Intersection over Union (mIoU) metric and instance segmentation with the mean Average Precision (mAP) metric.
@@ 5 Baseline Experiments
In this section, we describe various weak and strong baselines based on state-of-the-art architectures and methods for four FungiTastic benchmarks. We report results for the closed-set, few-shot learning, and zero-shot segmentation, but other baselines will be provided later in the supplementary materials, documentation, or on the dataset website.
@@ 5.1 Closed-set Image Classification
We train a variety of state-of-the-art CNN architectures to establish some baselines for closed-set classification on the FungiTastic and FungiTastic–M. All selected architectures were optimized with Stochastic Gradient Descent with momentum set to 0.9, SeeSaw loss [69], a mini-batch size of 64, and Random Augment [12] with a magnitude of 0.2. The initial LR was set to 0.01 (except for ResNet and ResNeXt, with LR=0.1), and it was scheduled based on validation loss.
Results: Similarly to other fine-grained benchmarks, while the number of params, complexity of the model, and training time are more or less the same, the transformer-based architectures achieved considerably better performance on both FungiTastic and FungiTastic–M and two different input sizes (see Table 3 and Table 8 in Appendix). The best-performing model, BEiT-Base/p16 [3], achieved $\text{F}_{1}^{m}$ just around 40%, which shows the severe difficulty.
Table 3: Closed-set fine-grained classification on FungiTastic and FungiTastic–M. A set of selected state-of-the-art Convolutional- (top section) and Transformer-based (bottom section) architectures evaluated on test sets. All reported metrics show the challenging nature of the dataset.
FungiTastic–M – 2242 FungiTastic – 2242
Architecture Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$
ResNet-50 [25] 61.7 79.3 35.2 62.4 77.3 32.8
ResNeXt-50 [71] 62.3 79.6 36.0 63.6 78.3 33.8
EfficientNet-B3 [62] 61.9 79.2 36.0 64.8 79.4 34.7
EfficientNet-v2-B3 [63] 65.5 82.1 38.1 66.0 80.0 36.0
ConvNeXt-Base [42] 66.9 84.0 41.0 67.1 81.3 36.4
ViT-Base/p16 [18] 68.0 84.9 39.9 69.7 82.8 38.6
Swin-Base/p4w12 [41] 69.2 85.0 42.2 69.3 82.5 38.2
BEiT-Base/p16 [3] 69.1 84.6 42.3 70.2 83.2 39.8
@@ 5.2 Few-shot Image Classification
Three baseline methods are implemented. The first baseline is standard classifier training with the Cross-Entropy (CE) loss. The other two baselines are nearest-neighbor classification and centroid prototype classification based on deep embeddings extracted from large-scale pre-trained vision models, namely CLIP [51], BioCLIP [59] and DINOv2 [45].
@@
Standard deep classifiers are trained with the CE loss to output the class probabilities for each input sample. Nearest neighbors classification ($\mathbf{k}$-NN) constructs a database of training image embeddings. At test time, $k$ nearest neighbors are retrieved, and the classification decision is made based on the majority class of the nearest neighbors. Nearest-centroid-prototype classification constructs a prototype embedding for each class by aggregating the training data embeddings of the given class. The classification depends on the image embedding similarity to the class prototypes. These methods are inspired by prototype networks proposed in [57].
@@
Results: While DINOv2 [45] embeddings greatly outperform CLIP [51] embeddings, BioCLIP [59] outperforms them both, highlighting the dominance of domain-specific models. Further, the centroid-prototype classification always outperforms the nearest-neighbor methods. Finally, the best standard classification models trained on the in-domain few-shot dataset underperform both DINOv2 and CLIP embeddings, which shows the power of methods tailored to the few-shot setup. For results summary, refer to Table 4.
@@
Table 4: Few shot classification on FungiTastic–Few-shot. Pretrained deep descriptors with the nearest centroid and 1-NN nearest neighbor classification (Left) and fully supervised (max 4 examples per class) classifier with cross-entropy-loss (Right). All pre-trained models are based on the ViT-B architecture, CLIP [51], and BioCLIP [59] with patch size 32 and DINOv2 [45] with patch size 16.
Model Method Top1 Top3
CLIP 1-NN 6.1
centroid 7.2 13.0
DINOv2 1-NN 17.4
centroid 17.9 27.8
BioCLIP 1-NN 18.8
centroid 21.8 32.6
Architecture Input Top1 Top3
BEiT-B/p16 224$\times$224 11.0 17.4
384$\times$384 11.4 18.4
ConvNeXt-B 224$\times$224 14.0 23.1
384$\times$384 15.4 23.6
ViT-Base/p16 224$\times$224 13.9 21.5
384$\times$384 19.5 29.0
5.3 Experiments with Additional Metadata
@@
We provide baseline experiments using tabular metadata (habitat, month, substrate) based on previous work [49]. Table LABEL:tab:exps:tab_data shows that all the attributes improve all the metrics. Individually, the addition of the habitat attribute results in the biggest gains in accuracy (2.3%), followed by substrate (1.2%) and month (0.9%). Overall, habitat was the most efficient way to improve performance. With the combination of Habitat, Substrate, and Month, we improved the EfficientNet-B3 model’s performance on FungiTastic-M by 3.62%, 3.42% and 7.46% in Top1, Top3, and F1, respectively, indicating the gains are mostly orthogonal. Using the MetaSubstrate instead of Substrate resulted in performance lower by 0.2%, 0.5%, and 0.3% in Top1, Top3, and F1, respectively.
@@
Table 5: Ablation on a combination of observation-related data. Utilizing a simple yet effective approach based on previous work [49], we measure performance improvement using Habitat, Substrate, and Month and their combination. We also test how replacing Substrate variables with MetaSubstrate affects performance. Evaluated with EfficientNet-B3 on FungiTastic-M test set.
@@ 5.4 Segmentation
Results: The baseline method achieved an average per-image IoU of 89.36%. While the model exhibits strong zero-shot performance, it sometimes fails to detect mushrooms. These instances often involve small mushrooms, where a higher input resolution could enhance detection, and atypical mushrooms, such as very thin ones. Another common issue is SAM’s tendency to miss mushroom stems. The results for the simplified foreground-background segmentation task underline the need for further development of domain-specific models. Qualitative results, including random images and examples where the segmentation performs best and worst, are reported in Figure 6.
@@
Figure 6: Zero-shot Fungi segmentations on FungiTastic–M benchmark. Random samples (top section), best IoU samples (mid), and worst IoU samples (bottom). Highlighted pixels correspond to: true positives, false positives, and false negatives.
5.5 Open-set Image Classification
While constructing baselines for the FungiTastic open-set benchmark, we approached open-set classification as a binary decision-making problem, where the model determines whether a new image belongs to a known class or a novel class. This method serves as an initial step in the classification pipeline, deciding if a closed-set classifier is suitable for recognizing a given sample. We evaluate several approaches for open-set classification:
  • Maximum Softmax Probability [28] (MSP): Uses the highest probability from softmax output from a closed-set classifier as the open-set score.
  • Maximum Logit Score [67] (MLS): Uses the highest logit output from a closed-set classifier as the open-set score.
  • Nearest Mean Score (NM): Computes the mean embedding for each class, then calculates the Euclidean distance between an image embedding and the nearest class mean.
We use features and logits from the BEiT-Base/p16 closed-set classifier baseline, trained on the full dataset (i.e., FungiTastic) at a 384$\times$384 resolution. To explore the potential of general pre-trained representation, we compare the fully-supervised model with generic features from a pre-trained DINOv2 model [45]. Using DINOv2 features, we train a simple linear layer to obtain MSP and MLS scores. Note that BEiT represents the best model from the closed-set classification baseline.
Results: The MLS method achieved the best open-set classification performance on both backbones. With BEiT-Base/p16, MLS achieves a TNR95 of 27.7% and an AUC of 83.9%, which are the highest AUC across all methods. DINOv2, in contrast, achieves the best TNR95 with a score of 36.9% using the MLS method, though its AUC is slightly lower at 74.5%. The MSP method also performs well with DINOv2, reaching a TNR95 of 32.5% and an AUC of 82.4%. However, the NM method, which relies on feature embeddings rather than classifier outputs, significantly underperforms in both metrics. See Table 6 for more details.
Table 6: Open-set classification baselines. Overall, the results are inconclusive and highly metric-dependent. The MLS (Max. Logit) method with the BEiT-Base/p16 backbone yields the highest AUC (83.9%), while the DINOv2 backbone with MLS achieves the highest TNR95 (36.9%). The NM (Nearest Mean) method consistently underperforms in both metrics across both backbones. For the AUC metric, MLS with a BEiT backbone (fine-tuned on the FungiTastic closed-set dataset) outperforms other approaches. However, both MSP (Max. Softmax) and MLS using DINOv2 linear layer are better when TNR95 performance is considered.
Nearest Mean Max. Logit Max. Softmax
Backbone TNR95 AUC TNR95 AUC TNR95 AUC
BEiT-Base/p16 23.2 73.9 27.7 83.9 25.3 79.8
DINOv2 12.1 69.2 36.9 74.5 32.5 82.4
5.6 Vision-Language Fusion
To evaluate the relevance of the available textual data (i.e., photograph captions) for species classification, we provide baselines that use a sequence classification variant of the lightweight DistilBERT [55] model trained as a classifier on textual descriptions only. The model was trained for 10 epochs using the standard cross-entropy loss, with logits obtained from a classification head applied to the pooled features of the class token in DistilBERT. For evaluation, we use text descriptions generated for the images in the test set.
Results: The DistilBERT classifier achieves a Top1 accuracy of 31.2% on FungiTastic-M and 24.1% on the full benchmark; significantly lower than fully supervised BEiT classifiers. However, this is still a strong result, given that it relies solely on textual descriptions. A simple ensemble that averages logits from the image and text classifiers shows potential for improved accuracy, indicating that the two methods are complementary. VLM-based descriptions capture useful details often missed by the image model. Further analysis shows the ensemble improves performance mainly on common categories, while the image classifier performs better on rare ones. This trade-off likely accounts for the drop in $\text{F}_{1}^{m}$ and overall accuracy on the full benchmark. For more details, see Table 7 and Figure 7.
Table 7: Vision-Language fusion performance. DistillBERT uses text descriptions of images for species classification. Fusion method predictions are the mean of DistillBERT and BEiT logits.
FungiTastic–M – 2242 FungiTastic – 2242
Architectures Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$
DistillBERT 31.2 50.2 11.5 24.1 39.1 8.8
BEiT-Base/p16 67.3 83.3 40.5 70.2 83.2 41.1
Fusion 67.7 83.8 39.8 69.0 82.6 40.0
Figure 7: Vision-Language fusion – accuracy dependence on class frequency. Like the vision model (BEiT-Base/p16), the language model (DistillBERT) struggles with infrequent classes. Fusion improves accuracy mainly for species with over 100 samples. The test set is binned by class frequency into deciles (x-axis).